OpenStudy (buttermequeasy):

Solve for x: −2|x − 3| = −12 x = −3, x = 4.5 x = 3, x = 9 x = −3, x = 9 No solutions

OpenStudy (buttermequeasy):

Please...

OpenStudy (anonymous):

−2|x − 3| = −12 given |x-3|=6 divide by -2 to isolate abs value x-3=-6 or x-3=6 definition of abs value x=-3 or x=9 add 3 to solve for x

OpenStudy (anonymous):

C. x = −3, x = 9